OpenStudy (anonymous):

Does anyone know how to do this one? Evaluate the line integral ∫ F • dr C about the simple closed curve C, oriented in the positive direction, where C is the circle x2 + y2 = 16 and F = (-4y)(sin2 x) i + (2x + sin 2x) j

OpenStudy (anonymous):

hold on i was just finished with line integrals =D

OpenStudy (anonymous):

use Green's theorem...

OpenStudy (anonymous):

...wait, if and when im done with this one, can you help me w/ a parametrization?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

so green's theorem says: doubleintegral(dQ/dx-dP/dy)dA

OpenStudy (anonymous):

so, dQ/dx=4x*cos(2x)+2sin(2x) dP/dy=-4sin(2x)

OpenStudy (anonymous):

doubleintegral(4x*cos(2x)-2sin(2x))dA

OpenStudy (anonymous):

ahh, wait you need to use a "power reducer"

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

cos(2x)=1-2sin^2(x) sin(2x)=2sin(x)cos(x) ...ugly

OpenStudy (anonymous):

can someone help me with my problem

OpenStudy (anonymous):

doubleintegral(4x-8x*sin^2(x)-4sin(x)cos(x))dA

OpenStudy (anonymous):

you with me?

OpenStudy (anonymous):

yeah i got you. But shouldnt we paramatize the x and y values

OpenStudy (anonymous):

wait, that's not going to work... hold on

OpenStudy (anonymous):

i guess you know more about this than i do, we didn't parametrize anything when we were doing these. yes, i figure you can do that

OpenStudy (anonymous):

I was thinking that b/c this was a circle, we'll do a change of variables into polar coordinates

OpenStudy (anonymous):

so..x=4cos(t) y=4sin(t) ?

OpenStudy (anonymous):

yeah, but right now, as far as I am aware, you can't integrate cos(cos(t))

OpenStudy (anonymous):

then t would be between 0 and 2pi..

OpenStudy (anonymous):

we have to work it out somehow so that the sines and cosines can disappear

OpenStudy (anonymous):

then we can do the change of variables

OpenStudy (anonymous):

so right now we are at doubleintegral(4x*cos(2x)-2sin(2x))dA

OpenStudy (anonymous):

oh wait I see.. using greens theorem M or P=-4(sin^2(x)) N or Q=2x+sin2x then, we derive with respect to x on N and derive with respect to y on M

OpenStudy (anonymous):

that's what we did

OpenStudy (anonymous):

the problem is the double integral

OpenStudy (anonymous):

seems that i made a mistake when i did dQ/dx it's supposed to be 2cos(2x)+2

OpenStudy (anonymous):

what would be the bounds of integration then?

OpenStudy (anonymous):

right now, we haven't "parametrized" in polar coordinates, the integral would be from 0 to 2pi, 0 to 4

OpenStudy (anonymous):

ok..i see

OpenStudy (anonymous):

it's actually doubleintegral(4x*cos(2x)+2sin(2x)+2)dA

OpenStudy (anonymous):

i mean... doubleintegral(2cos(2x)+4sin(2x)+2)dA oops, can't keep track

OpenStudy (anonymous):

know of any property that can reduce this so that the trig disappears?

OpenStudy (anonymous):

hmm..well is there an identity for cos(2x) or sin(2x) ?

OpenStudy (anonymous):

yes, though i don't think it's going to help.

OpenStudy (anonymous):

you might try a classic integration, with y from -sqrt(4-x^2) to sqrt(4-x^2) and x from -4 to 4

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

in that case you need a calculator to help you with arcsin and what not

OpenStudy (anonymous):

im doing that now, but the calc's going terribly slowly

OpenStudy (anonymous):

this is not going to work, is it?

OpenStudy (anonymous):

i dont know I got a 32sqrt(4-x^2) using wolfram mathmatics

OpenStudy (anonymous):

see, now if we did the problem without using green's theorem, it won't work anyway, since F(r(t)) gets us another cosine within a cosine.

OpenStudy (anonymous):

remember, it's a double integral, x should be eliminated

OpenStudy (anonymous):

wait a minute, that sin2x you wrote, it that a "sine squared x" or a "sine double x"???

OpenStudy (anonymous):

sin squared

OpenStudy (anonymous):

darn, you caused me soooo much hell.

OpenStudy (anonymous):

this is much easier than i thought

OpenStudy (anonymous):

I mean no harm. I come in peace.

OpenStudy (anonymous):

jk =D

OpenStudy (anonymous):

can you rewrite the question a bit neater so i know what exactly im integrating?

OpenStudy (anonymous):

"sine squared x" should be (sin(x))^2

OpenStudy (anonymous):

ok..(-4y)(sin(x))^2 i + (2x + sin 2x) j

OpenStudy (anonymous):

hmmm...

OpenStudy (anonymous):

and 2x + sin2x i presume is "sine double x"?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

much better

OpenStudy (anonymous):

lets start over

OpenStudy (anonymous):

dQ/dx=2cos(2x)+2 dP/dy=-4(sin(x))^2

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

did i get the derivative wrong?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

perfect works out very nicely

OpenStudy (anonymous):

so doubleintegral(dQ/dx-dP/dy)dA is doubleintegral(2cos(2x)+2+4(sin(x))^2)dA

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

using cos(2x)=2(cos(x))^2-1 we get doubleintegral(2(2cos(x))^2+2+4(sin(x))^2)dA

OpenStudy (anonymous):

which is doubleintegral(4(cos(x))^2-2+2+4(sin(x))^2)dA which is doubleintegral(4(cos(x))^2+4(sin(x))^2)dA

OpenStudy (anonymous):

follow?

OpenStudy (anonymous):

yesm

OpenStudy (anonymous):

ok, oops, typo in doubleintegral(2(2cos(x))^2+2+4(sin(x))^2)dA: should be doubleintegral(2(2cos(x)-1)^2+2+4(sin(x))^2)dA

OpenStudy (anonymous):

arrgh, i mean doubleintegral(2((2cos(x))^2-1)+2+4(sin(x))^2)dA

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

you understand, though? didn't mean to have all those typos in the way.

OpenStudy (anonymous):

that s what got us into this mess ;)

OpenStudy (anonymous):

ok so now we have doubleintegral(4(cos(x))^2+4(sin(x))^2)dA easy cake: (cos(x))^2+(sin(x))^2=1, so, it is just doubleintegral(4)dA

OpenStudy (anonymous):

now, you can do a change of variables!

OpenStudy (anonymous):

doubleintegral[theta,0,2pi][r,0,4](4r)dr dtheta evaluate that, and you're done

OpenStudy (anonymous):

do you get 32pi when you integrate r from 0 to 4 and theta from 0 to to pi

OpenStudy (anonymous):

64pi

OpenStudy (anonymous):

im pretty sure

OpenStudy (anonymous):

well, no r integrated is 16 pi, the full solution i worked out to be 64 pi

OpenStudy (anonymous):

you were supposed to integrate r from 0 to 4 and theta from 0 to 2pi, not pi.

OpenStudy (anonymous):

thats what i meant 2pi i still get 32pi

OpenStudy (anonymous):

integrate r, which is r^2/2. 4^2/2-0=8 integrate 1, which is theta 2pi-0=2pi multiply you get 16 pi

OpenStudy (anonymous):

oooooh....ok

OpenStudy (anonymous):

you are oh so smart

OpenStudy (anonymous):

nahhh, it was nothing i got problems of my own.

OpenStudy (anonymous):

i'll post my question (for the 4th time) =D